# Segmenting Action-Value Functions Over Time-Scales in SARSA via TD($\Delta$)

### v2

@@ Abstract
In numerous episodic reinforcement learning (RL) settings, SARSA-based methodologies are employed to enhance policies aimed at maximizing returns over long horizons. Conventional SARSA algorithms, however, have difficulties in balancing bias and variation due to the reliance on a singular, fixed discount factor. This study expands the temporal difference decomposition approach, TD($\Delta$), to the SARSA algorithm, which we designate as SARSA($\Delta$). SARSA, a widely utilised on-policy RL method, enhances action-value functions via temporal difference updates. TD($\Delta$) facilitates learning over several time-scales by breaking the action-value function into components associated with distinct discount factors. This decomposition improves learning efficiency and stability, particularly in problems necessitating long-horizon optimization. We illustrate that our methodology mitigates bias in SARSA’s updates while facilitating accelerated convergence in both deterministic and stochastic environments. Experimental findings across many benchmark tasks indicate that the proposed SARSA($\Delta$) surpasses conventional TD learning methods in both tabular and deep RL environments.
@@
SARSA, Temporal Difference, Action-Value Functions, On-policy, Finite Horizon
@@ 1 Introduction
In reinforcement learning (RL), the objective is to enhance an agent’s actions to maximize cumulative rewards through interaction with an environment. Temporal difference (TD) learning methods, like Q-learning and SARSA, have demonstrated efficacy across several tasks by allowing agents to estimate action-value functions that forecast anticipated future rewards. Traditionally, these strategies utilise a discount factor $10\leq\eta<$ 1, with values approaching $\eta=0$ to emphasise short-term rewards over long-term ones, therefore diminishing the planning horizon and enhancing learning stability and efficiency[1]. Discount factors $\eta<1$ frequently yield enhanced results during the early phases of learning, as demonstrated by Prokhorov and Wunsch[2]. In specific contexts, such as long-horizon tasks[3, 4] where long-term planning is essential—like navigation or decision-making tasks with deferred rewards—this method can create bias, constraining the agent’s capacity to acquire optimal long-horizon policies.
For instance: SARSA[1] (State-Action-Reward-State-Action) is an on-policy RL algorithm that modifies an action-value function Q(s,a) according to the state-action transitions encountered by the agent. The conventional SARSA algorithm performs well in contexts with shorter time horizons; however, in long-horizon tasks, the singular discount factor $\eta$ employed in SARSA results in a compromise between bias and variance. Small values of $\eta$ yield myopic policies, whereas larger values induce substantial variance in the estimated action-value functions, rendering the learning process unstable and inefficient. In response, other solutions have been suggested, including[4, 2, 5, 6, 7, 8].
Recent research in TD learning has investigated the segmentation of value functions across multiple time scales, enabling agents to enhance value estimates concurrently at different horizons[8]. The TD($\Delta$) architecture improves scalability and performance in scenarios involving long-term dependencies by breaking down value functions into components that incorporate different discount factors, known as TD($\Delta$). The adaptability of the TD($\Delta$) framework shows considerable promise in addressing the bias-variance trade-offs that are characteristic of TD learning.
@@
We propose SARSA($\Delta$), an enhancement of the TD($\Delta$) approach integrated into the SARSA algorithm, based on this principle. TD($\Delta$) emphasizes the actor-critic framework and conventional TD learning, whereas SARSA($\Delta$) divides the state-action value function into multiple partial estimators corresponding to various discount factors, known as delta estimators. These estimators approximate the difference $D_{m}(s,a)=Q_{\eta_{m}}(s,a)-Q_{\eta_{m-1}}(s,a)$ between action-value functions, thereby improving learning efficiency in scenarios where actions affect long-horizon outcomes. Similar to TD($\Delta$), SARSA($\Delta$) seeks to optimize a set of delta estimators, where each estimator is associated with a distinct discount factor. With smaller discount factors, convergence happens faster because to segmentation, while greater discount factors strengthen this foundation for better long-horizon planning.
This study introduces a multi-step SARSA($\Delta$) update rule that disaggregates action-value functions, facilitating accelerated convergence and enhanced policy efficacy, especially in intricate deterministic and stochastic environments. Furthermore, we illustrate that SARSA($\Delta$) may seamlessly adapt to any multi-step reinforcement learning techniques, including n-step SARSA and eligibility traces. Through theoretical analysis and empirical findings, we emphasise the benefits of SARSA($\Delta$) in diverse tasks, including a basic ring MDP utilised by Kearns & Singh[9], demonstrating its superiority over conventional SARSA in scenarios necessitating a balance between short-term action consequences and long-term reward maximisation. SARSA($\Delta$) possesses the capability to adapt discount factors across varying time scales, serving as an effective instrument in environments where time scale is critical.
@@
The impetus for extending TD($\Delta$) to SARSA($\Delta$) lies in the fact that temporal difference (TD($\Delta$)) learning offers a framework for deconstructing value functions across many time scales, hence overcoming the constraints of single-scale learning in long-horizon problems. By extending TD($\Delta$) to SARSA($\Delta$), we seek to decompose the action-value function Q(s,a) into elements that emphasise distinct discount factors. This decomposition enables SARSA to utilise the advantages of learning across several time scales, hence stabilising the learning process and mitigating the bias-variance tradeoff. This addon provides a more detailed and scalable method for action-value function learning in intricate environments.
The primary contributions of this paper are as follows: (i) We expand TD($\Delta$) to SARSA($\Delta$) by partitioning the action-value function Q(s,a) into delta components over several time-scales. (ii) The suggested SARSA($\Delta$) algorithm enhances learning efficiency and stability by independently learning components for various discount factors and integrating them to represent the complete action-value function. (iii) Both theoretical and empirical evaluations have shown that this multi-scale breakdown has a number of benefits, particularly in environments where there are long-horizon rewards.
@@
The ensuing portions of the paper are organized as follows: Section 2 presents a discourse on pertinent literature. Section 3 provides the essential background information. The principles of SARSA($\Delta$) and theoretical analysis are shown in Sections 4 and 5, respectively. Section 6 provides a detailed summary of the experiments performed. Ultimately, we present a conclusion in section 7.
@@ 2 Related Work
In reinforcement learning (RL), optimising for long-horizon rewards is tough due to the complexity of learning with an undiscounted return. Temporal discounting is frequently employed to streamline this process; nevertheless, it may induce bias. Researchers[8] have investigated techniques to separate value functions across various time scales by analysing disparities among value functions with lower discount factors, hence enhancing scalability and performance. The study introduces $\gamma$-nets[10], a technique for generalising value function estimate across timescales in RL. The authors of the paper[11] present an approach for integrating hyperbolic discounting into RL. Traditional RL employs exponential discounting, which does not align with the hyperbolic discounting observed in human and animal behaviors. The authors present an agent designed to approximate hyperbolic discount functions by utilizing temporal-difference learning methods. Furthermore, the study reveals that learning value functions across multiple time horizons improves performance, especially when implemented as an auxiliary task in value-based RL algorithms such as Rainbow. This method demonstrates potential in contexts characterized by uncertainty and risk. Expected SARSA[7], a variant of SARSA, utilizes information regarding the stochastic nature of the behavior policy to execute updates with reduced variance, which speeds up learning. This study[12] presents temporal-difference (TD) learning methods, which underpin numerous RL algorithms, notably SARSA, by forecasting future outcomes based on the discrepancies between consecutive forecasts rather than the ultimate result. The research demonstrates the convergence of TD techniques under specific situations and emphasises their benefits in predictive tasks, particularly when incremental learning is crucial. This study[13] presents theoretical convergence results for temporal-difference learning methods utilising function approximation. Extending this analysis to SARSA($\Delta$) enables us to assert comparable convergence properties, provided that the delta components are updated in accordance with the SARSA update rule. It is noteworthy that none of the previously mentioned works employ short-term estimates for training long-term action-value functions. Our method can also be utilised as a generalised action-value function, and its capacity to query smaller time-scales represents an additional advantage. Performance improvements primarily arise from the separation of action-value functions across varying time scales through the use of TD($\Delta$) in SARSA.
Recently, there has been some research conducted on the precise selection of the discount factor[2, 5, 6, 4]. The authors[6] of this study introduce meta-gradient methods for the adaptive selection of discount factors in RL, emphasising the necessity of balancing short- and long-term rewards, which corresponds with the rationale for dissecting action-value functions across several discount factors in SARSA($\Delta$).
Lastly, another extensive category of study pertinent to our research indirectly encompasses hierarchical RL and the bias-variance trade-off. These studies[14, 15, 16, 17, 18, 19, 20] present the concept of deconstructing value functions into subtasks by hierarchical RL, analogous to the manner in which TD($\Delta$) disaggregates value functions across temporal scales. This substantiates the assertion that value function decomposition can enhance learning efficacy and stability. This paper[9] functions as a theoretical foundation for comprehending the bias-variance trade-off that arises when decomposing Q-values across various time-scales, and it also establishes constraints on the bias and variance of TD updates. This study[21] presents empirical findings demonstrating the efficacy of deep RL techniques utilizing value function approximation in Atari-like environments. It also endorses the necessity for techniques that diminish variance in long-horizon tasks.
@@ 3 Background and notation
Examine a completely observable Markov Decision Process (MDP) [22], characterized by the tuple $(\mathcal{S},\mathcal{A},\mathcal{P},r)$, where $\mathcal{S}$ denotes the state space, $\mathcal{A}$ signifies the action space, and $\mathcal{P}:\mathcal{S}\times\mathcal{A}\rightarrow\mathcal{S}\rightarrow[0,1]$ represents the transition probabilities that associate state-action pairings with distributions across subsequent states, whereas $r:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$ denotes the reward function. At each timestep $n$, the agent occupies state $s_{n}$, chooses an action $a_{n}$, obtains a reward $r_{n}=r(s_{n},a_{n})$, and transitions to the subsequent state $s_{n+1}\sim\mathcal{P}(\cdot|s_{n},a_{n})$.
In a conventional Markov Decision Process (MDP) framework, an agent seeks to maximize the discounted return defined as $Q_{\eta}^{\pi}(s,a)=\mathbb{E}\left[\sum_{n=0}^{\infty}\eta^{n}r_{n+1}\mid s_{n}=s,a_{n}=a\right]$, where $\eta$ denotes the discount factor and $\pi:\mathcal{S}\rightarrow\mathcal{A}\rightarrow[0,1]$ represents the policy adhered to by the agent. The action-value function $Q_{\eta}^{\pi}(s,a)$ is determined as the fixed point of the Bellman operator $\mathcal{T}Q=r^{\pi}+\eta\mathcal{P}^{\pi}Q$, where $r^{\pi}$ denotes the expected immediate reward and $\mathcal{P}^{\pi}$ represents the transition probability operator associated with the policy $\pi$. For convenience, we remove the superscript $\pi$ for the rest of the paper.
@@
An on-policy RL technique called SARSA learns the action-value function Q(s,a) for a given policy $\pi$. The action-value function represents the expected total reward that comes from state s, taking action a, and following policy $\pi$. Then the SARSA update rule is clarified as follows:
@@
On the other hand, for a complete trajectory, we can employ the discounted sum of one-step TD errors, generally known as the $\lambda$-return[23] or, equivalently, the Generalized Advantage Estimator (GAE)[24]. The GAE enhances advantage estimates by reconciling the trade-off between variance and bias via parameters $\lambda$ and $\eta$. The equation for the Generalized Advantage Estimator is generally expressed as:
@@
Where $\delta_{n+k}$ is the TD error at time n, computed as follows: $\delta_{n+k}=r_{n+k}+\eta Q(s_{n+k+1},a_{n+k+1})-Q(s_{n+k},a_{n+k})$
@@
Proximal Policy Optimization (PPO) [28], an enhancement of actor-critic methodologies, confines policy updates to a designated optimization area, referred to as a trust region, utilizing a clipping objective that evaluates the current parameters, $\nu$, against the prior parameters, $\nu_{old}$:
@@
In on-policy methods like SARSA, the likelihood ratio $\rho(\nu)$ between the current policy and a prior policy is typically employed. SARSA evaluates actions in accordance with the current policy, concurrently updating the action-value function as actions are performed. However, when utilized within an actor-critic framework, the SARSA-learning agent may adopt a similar policy ratio:

### v4

@@ Abstract
In numerous episodic reinforcement learning (RL) environments, SARSA-based methodologies are employed to enhance policies aimed at maximizing returns over long horizons. Traditional SARSA algorithms face challenges in achieving an optimal balance between bias and variation, primarily due to their dependence on a single, constant discount factor ($\eta$). This investigation enhances the temporal difference decomposition method, TD($\Delta$), by applying it to the SARSA algorithm, now designated as SARSA($\Delta$). SARSA is a widely used on-policy RL method that enhances action-value functions via temporal difference updates. By splitting the action-value function down into components that are linked to specific discount factors, SARSA($\Delta$) makes learning easier across a range of time scales. This analysis makes learning more effective and ensures consistency, particularly in situations where long-horizon improvement is needed. The results of this research show that the suggested strategy works to lower bias in SARSA’s updates and speed up convergence in both deterministic and stochastic settings, even in dense reward Atari environments. Experimental results from a variety of benchmark settings show that the proposed SARSA($\Delta$) outperforms existing TD learning techniques in both tabular and deep RL environments.
@@
SARSA, Temporal Difference (TD($\Delta$)), Action-Value Functions, On-policy, Long-Horizon, Time-Scale Decomposition
@@ 1 Introduction
Reinforcement learning (RL) agents encounter significant challenges in the optimization of long-horizon rewards. Traditional methods employing singular discount variables encounter inherent challenges: small $\eta$ results in low-variance but myopic policies, whereas large $\eta$ suffers from excessive variance. Fixed discounts remain constant despite variations in reward densities[1, 2, 3]. Temporal difference (TD) learning methods, such as Q-learning and SARSA, have shown effectiveness in various tasks by enabling agents to estimate action-value functions that predict expected future rewards. Traditionally, these strategies use a discount factor of $0\leq\eta<1$, with values getting closer to $\eta=0$ to put more weight on short-term rewards than long-term ones. This approach shortens the planning horizon and makes learning more stable and efficient[2]. Prokhorov and Wunsch[4] show that discount factors $\eta<1$ often lead to better results in the early stages of learning. In certain scenarios, such as long-horizon tasks[5, 6] where long-term planning is crucial—like navigation or decision-making tasks that involve delayed rewards—this approach can introduce bias, complicating the agent’s ability to determine the most effective long-horizon policies.
For example, SARSA[2] (State-Action-Reward-State-Action) is an on-policy RL algorithm that updates the action-value function Q(s,a) according to the state-action transitions the agent sees. The standard SARSA algorithm works well in environments with shorter time horizons; however, it faces challenges when utilized for tasks with longer horizons. The main reason for the observed phenomenon is the use of a single discount factor $\eta$, which requires a trade-off between bias and variance. This often leads to instability and inefficiency during the learning process. A variety of recommendations exist for addressing this issue [6, 4, 7, 8, 9, 1, 10].
Through the use of weighted combinations of value functions, Fedus et al. [10] presented the concept of hyperbolic discounting. However, their formulation is restricted to state-value functions and does not extend to action-value functions. Meta-gradient RL[8] is a technique that enables agents to adapt their learning processes dynamically. Meta-gradient means finding or improving targets and hyperparameters during training, which could make learning more efficient and adaptable. A gradient-based meta-learning algorithm capable of online adaptation to the true value function while engaging with and learning from the environment. However, akin to other methodologies, meta-gradient RL primarily emphasizes state-value functions, rather than action-value formulations.
The work that is most relevant in this case is by Romoff et al.[1], which provides a theoretical basis for temporal difference decomposition using the TD($\Delta$) framework. TD($\Delta$) segments the state-value function V(s) into components using different discount factors, also known as TD($\Delta$). This lets the agent learn about short-term and long-term returns separately, which can then be combined to improve performance. This decomposition framework demonstrates considerable potential in addressing the bias-variance trade-off that is intrinsic to temporal difference learning. However, TD($\Delta$) is constrained to state-value functions, which presents a limitation in the application of multi-timescale representations to action-value-based algorithms such as SARSA and Q-learning.
@@
This article proposes SARSA($\Delta$), an enhancement of the TD($\Delta$) approach integrated into the SARSA algorithm, based on this principle. TD($\Delta$) emphasizes the actor-critic framework and conventional TD learning, whereas SARSA($\Delta$) divides the state-action value function into multiple partial estimators corresponding to various discount factors, known as delta estimators. These estimators give a rough estimate of the difference $D_{m}(s,a)=Q_{\eta_{m}}(s,a)-Q_{\eta_{m-1}}(s,a)$ between action-value functions. This makes learning faster in contexts where actions yield long-term effects. Like TD($\Delta$), SARSA($\Delta$) aims to enhance a collection of delta estimators, with each estimator linked to a specific discount factor. Smaller discount factors lead to quicker convergence via segmentation, whereas larger discount factors enhance this foundation, facilitating improved long-term planning.
This paper provides a multi-stage SARSA($\Delta$) updating algorithm that separates action-value functions, which allows for faster convergence and improved policy efficacy, especially in complex deterministic and stochastic environments like Atari games. Moreover, our research shows that SARSA ($\Delta$) can seamlessly accommodate various multi-step RL methodologies, such as n-step SARSA and eligibility traces. This study emphasizes the benefits of SARSA($\Delta$) in various tasks, including a basic ring MDP used by Kearns & Singh[11]. It outperforms conventional SARSA in scenarios requiring a balance between short-term action consequences and long-term reward maximization. SARSA ($\Delta$) may adjust discount factors for multiple time scales, making it valuable where time scale is precious.
@@
The reason for extending TD($\Delta$) to SARSA($\Delta$) is that temporal difference (TD($\Delta$)) learning provides a way to break down value functions across a range of time scales, which solves the problems that come up when learning on a single scale for long-horizon scenarios. We want to break down the action-value function Q(s,a) into components that focus on different discount factors by extending TD($\Delta$) to SARSA($\Delta$). This decomposition enables SARSA to leverage the advantages of learning over multiple time scales, which stabilizes the learning process and reduces the bias-variance trade-off. This add-on gives you a more detailed and scalable way to learn the action-value function in complex environments. Table 1 presents a comparison of the key features of TD($\Delta$)/SARSA($\Delta$) and Standard TD/SARSA-Learning.
Table 1: Comparison of the essential properties of TD($\Delta$) and SARSA($\Delta$).
Feature TD($\Delta$)[1] SARSA($\Delta$)-Learning (proposed by us) Standard TD[12] SARSA-Learning[12]
Value function decomposition Yes (by $\eta$) No No No
Action-value function decomposition No Yes (by $\eta$) No No
Handles multiple time-scales Yes Yes No No
Learning stability Enhanced, particularly long-term Enhanced, particularly long-term Reduced for long-term rewards Reduced for long-term rewards
Scalability Improved Improved Restricted by variance/bias Restricted by variance/bias
Recent research on value decomposition[1] has demonstrated potential for state-value functions; however, it identifies three significant shortcomings that this work aims to address: (i) Action-Value Formulation: No current method effectively decomposes Q(s, a) while maintaining on-policy guarantees. This study extends TD($\Delta$) to SARSA($\Delta$) by decomposing the action-value function Q(s,a) into delta components across multiple time scales. (ii) Theoretical Foundations: Current analyses do not address the bias-variance trade-off concerning action values. (iii) The proposed SARSA($\Delta$) algorithm improves learning efficiency and stability through the independent learning of components corresponding to different discount factors, which are subsequently integrated to form the comprehensive action-value function. (iv) Both theoretical and empirical evaluations have shown that this multi-scale breakdown has a number of benefits, particularly in environments where there are long-horizon rewards, as demonstrated in Table 1.
@@
The subsequent portions of the paper are organized as follows: Section 2 presents a discussion on relevant literature. Section 3 provides the essential background information. The principles of SARSA($\Delta$) and theoretical analysis are shown in Sections 4 and 5, respectively. Section 6 provides a detailed summary of the experiments performed. Ultimately, we present a conclusion in section 7.
@@ 2 Related Work
Optimizing for long-horizon rewards in reinforcement learning (RL) is difficult owing to the complexities of learning with undiscounted returns. Temporal discounting is often used to simplify this process; however, it can introduce bias.
2.1 Temporal Decomposition Approaches
In their investigation, Romoff et al. [1] explored the differentiation of value functions across multiple time scales by scrutinizing individuals with diminished discount factors. The system became more efficient and simpler to expand. In RL, Sherstan et al.[13] introduced $\gamma$-nets, a method designed to enhance value function estimates across various timescales. Authors’ recent work [10, 14] showed how to add hyperbolic discounting to RL. Exponential discounting was used in traditional RL, but this didn’t match the hyperbolic discounting seen in the behavior of human and animals. The authors showed an agent that used temporal-difference learning methods to get close to hyperbolic discount functions. The research also found that learning value functions across multiple time horizons increases performance, particularly when used as an auxiliary task in value-based RL algorithms like Rainbow. This approach has shown potential in risky and uncertain environments.
Recent research [15, 4, 7, 8, 6, 16, 17] focuses on the exact selection of the discount factor. Xu 124 et al. [8] proposed meta-gradient methods for choosing discount factors in RL that change over time. They stressed the importance of balancing short- and long-term rewards, which is similar to why we break down action-value functions across multiple discount factors in SARSA($\Delta$). Lastly, hierarchical RL and the bias-variance trade-off are part of a large body of research that is related to our study in an indirect way. These studies [18, 19, 20, 21, 22, 23, 24] introduced the idea of hierarchical RL breaking value functions down into smaller tasks, similar to how TD($\Delta$) broke value functions down across different time scales. These results backed up the claim that breaking down value functions can make learning more effective and stable.
2.2 Multi-Scale RL
To address issues in RL, especially when the optimum value function is complex and poorly represented by conventional DRL methods, van Seijen et al. [24] introduced Hybrid Reward Architectures (HRA). The HRA was designed to enhance learning consistency and effectiveness by partitioning the reward function and allocating distinct value functions to each component. Efficient multi-horizon learning in off-policy RL aims to improve agents’ ability to make predictions and decisions over a wide range of time scales. This is important for navigating through complex and unpredictable environments. Ali 141 et al. [25, 26] investigated designs that enable agents to learn value estimates over several horizons at the same time. This improves their capacity to adjust and flourish in novel environments. In the realm of artificial intelligence and robotics, coordinating agents with limited communication and across diverse planning periods is a significant challenge to multi-horizon, multi-agent planning.
Seiler et al. [27] introduced Decentralized Monte Carlo Tree Search (Dec-MCTS) and its variants, which have arisen as formidable instruments for addressing these challenges, allowing agents to plan effectively and adaptively in dynamic, unpredictable environments. Benechehab et al.[28] came up with a novel technique in model-based RL (MBRL) called multi-timestep models, aimed at mitigating the problem of escalating prediction errors over extended simulated trajectories. By optimizing for several future steps instead of only one-step-ahead forecasts, these models enhance the precision and resilience of long-term predictions, especially in complex or chaotic environments.
Bonnet et al. [29] presented multi-step meta-gradient RL, which uses meta-gradients that have been built up across many stages instead of just one. This makes learning algorithms more flexible. This method may provide richer and more useful learning signals, but it does come with some significant trade-offs. But it also increases computational complexity and variation, which may hinder performance if not handled carefully. While deep RL (DRL) can teach agents difficult tasks, it still struggles to improve sampling efficiency and adapt to changing environments. Eligibility traces, a well-established RL approach, expedite learning, but parameter dependencies make their integration with deep neural networks difficult. Kobayashi[30] introduced several time-scale eligibility traces to equilibrate short-term and long-term credit assignment. This technique improves learning speed and policy quality in online DRL environments by substituting the most important adaptively gathered traces.
2.3 SARSA Variants and Traditional RL Methods
Expected SARSA [9] is a kind of SARSA that speeds up learning by exploiting information from a stochastic behavior policy to create updates that are less variance. Many RL algorithms, like SARSA, employ temporal-difference (TD) learning, which looks at the differences between two predictions made one after the other instead of the final result. Sutton[31] came up with this notion for the first time in 1988. The research shows that TD approaches may be useful for predicting jobs, especially where incremental learning is crucial, and that they can even operate jointly in certain cases.
Theoretical convergence for function approximation-based TD learning algorithms was shown by Tsitsiklis and Van Roy[32]. If you follow the SARSA update rule for delta component updates, this research may be used with SARSA($\Delta$) to get the same convergence properties. None of the previously mentioned research discussed short-term estimations to train long-term action-value functions. One benefit is that you may ask about shorter time periods and utilize our method as a generic action-value function. Separating action-value functions across time scales using TD($\Delta$) in SARSA makes a big difference in performance. Theoretical work by Kearns and Singh[11] helped us comprehend the bias-variance trade-off that comes up when we break down Q-values over multiple time scales. It also set limits on the bias and variance of TD updates.
Most of the methods listed above focus on state values, although action-value decomposition has not yet been looked at. Our proposed strategy focuses on breaking down $Q(s,a)$ over different time scales, which makes long-term learning more stable.
@@ 3 Background and notation
Examine a completely observable Markov Decision Process (MDP) [33], characterized by the tuple $(\mathcal{S},\mathcal{A},\mathcal{P},r)$, where $\mathcal{S}$ denotes the state space, $\mathcal{A}$ signifies the action space, and $\mathcal{P}:\mathcal{S}\times\mathcal{A}\rightarrow\mathcal{S}\rightarrow[0,1]$ represents the transition probabilities that associate state-action pairings with distributions across subsequent states, whereas $r:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$ denotes the reward function. At each timestep $n$, the agent is state $s_{n}$, selects an action $a_{n}$, receives a reward $r_{n}=r(s_{n},a_{n})$, and transitions to the next state $s_{n+1}\sim\mathcal{P}(\cdot|s_{n},a_{n})$.
Within a standard MDP framework, an agent tries to get the highest possible discounted return, which is defined as $Q_{\eta}^{\pi}(s,a)=\mathbb{E}\left[\sum_{n=0}^{\infty}\eta^{n}r_{n+1}\mid s_{n}=s,a_{n}=a\right]$, where $\eta$ is the discount factor and $\pi:\mathcal{S}\rightarrow\mathcal{A}\rightarrow[0,1]$ stands for the policy that the agent follows. The action-value function $Q_{\eta}^{\pi}(s,a)$ is determined as the fixed point of the Bellman operator $\mathcal{T}Q=r^{\pi}+\eta\mathcal{P}^{\pi}Q$, where $r^{\pi}$ denotes the expected immediate reward and $\mathcal{P}^{\pi}$ represents the transition probability operator associated with the policy $\pi$. For convenience, we remove the superscript $\pi$ for the rest of the paper.
@@
An on-policy RL technique called SARSA learns the action-value function Q(s, a) for a given policy $\pi$. The action-value function represents the expected total reward derived from state s, executing action a, and adhering to policy $\pi$. Then the SARSA update rule is expressed as follows:
@@
On the other hand, for a complete trajectory, we can employ the discounted sum of one-step TD errors, generally known as the $\lambda$-return [34] or, equivalently, the Generalized Advantage Estimator (GAE) [35]. The GAE improves advantage estimates by balancing the trade-off between variance and bias through the parameters $\lambda$ and $\eta$. The Generalized Advantage Estimator is typically represented by the following equation:
@@
Where $\delta_{n+k}$ is the TD error at time n, computed as follows: $\delta_{n+k}=r_{n+k}+\eta Q(s_{n+k+1},a_{n+k+1})-Q(s_{n+k},a_{n+k})$.
@@
Proximal Policy Optimization (PPO) [39] is an improvement on actor-critic methods. It limits policy updates to a specific area of optimization called a trust region. It does this by using a clipping objective that compares the current parameters, $\nu$, to the old parameters, $\nu_{old}$:
@@
In on-policy approaches such as SARSA, the probability ratio $\rho(\nu)$ between the current policy and a preceding policy is often used. SARSA evaluates actions in accordance with the current policy, concurrently updating the action-value function as actions are performed. However, when utilized within an actor-critic framework, the SARSA-learning agent may adopt a similar policy ratio: